OpenStudy (anonymous):

Determine the standard form of the equation of the line that passes through (9, -4) and (6, 4) Haaaaalp

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi!

OpenStudy (misty1212):

first we need the slope, because no matter what, if you want the equation of a line you need to have the slope do you know how to find ti?

OpenStudy (anonymous):

that would be -4+4 over 9+4 right?

OpenStudy (misty1212):

not quite

OpenStudy (misty1212):

you want to subtract, not add

OpenStudy (anonymous):

Oh, okay. Hold on Lemme try.

OpenStudy (misty1212):

kk

OpenStudy (anonymous):

ok so its -8/5 ??

OpenStudy (misty1212):

hmm \(9-6=?\)

OpenStudy (anonymous):

oh dang! I was doing -4. My bad. So it's 4.

OpenStudy (misty1212):

lets go slow

OpenStudy (anonymous):

Yes

OpenStudy (misty1212):

(9, -4) and (6, 4)

OpenStudy (misty1212):

first we do it the think way, then we do it the math teacher way

OpenStudy (anonymous):

Okay, it's like the third day of school. My brain doesn't work yet! sorry ahah.

OpenStudy (misty1212):

from 6 to 9 in the x is right 3 units from 4 to -4 in the y is down 8 units right 3, down 8 slope is \[-\frac{8}{3}\]

OpenStudy (anonymous):

okay, yes.

OpenStudy (misty1212):

now the math teacher way \[\frac{y_2-y_1}{x_2-*x_1}=\frac{-4-4}{9-6}=-\frac{8}{3}\] or \[\frac{4-(-4)}{6-9}=\frac{8}{-3}=-\frac{8}{3}\] either way

OpenStudy (anonymous):

okay, got it.

OpenStudy (misty1212):

who goes to school in the middle of august?

OpenStudy (misty1212):

ok no matter, now we have the the slope, and we can use the"point slope" formula because we have two points to choose form

OpenStudy (anonymous):

Everyone where I'm from starts in August. I know some places in California start later. so y=mx+b?

OpenStudy (anonymous):

wait, thats not right.

OpenStudy (misty1212):

no that is the "slope intercept " form we are not there yet

OpenStudy (anonymous):

y-y1 = m (x-x1)

OpenStudy (misty1212):

yeah that one

OpenStudy (anonymous):

Okay so -4 - 4 = -8/3 (9 - 6) ?

OpenStudy (misty1212):

ooh no

OpenStudy (misty1212):

the x and y in the formula are the variables you see in the answer only the \(x_1\) and \(y_1\) are numbers (as well as m )

OpenStudy (misty1212):

you put numbers everywhere

OpenStudy (anonymous):

oh. gosh ok. y+4= -8/3 (x-9)

OpenStudy (misty1212):

looks a lot better now

OpenStudy (anonymous):

subtracting a negative turns it into an addition basically, right?

OpenStudy (misty1212):

right

OpenStudy (misty1212):

now you should probably put it in "slope intercept" form aka solve for y

OpenStudy (misty1212):

you know how to do that? takes the same two steps every time

OpenStudy (anonymous):

Move the 4 to the other side?

OpenStudy (misty1212):

no not yet

OpenStudy (misty1212):

first distribute the \(-\frac{8}{3}\) on the left

OpenStudy (anonymous):

oh yeah. so I multiply it or add it to the left?

OpenStudy (misty1212):

multiply

OpenStudy (misty1212):

\[y+4=-\frac{8}{3}+24\] if my arithmetic is correct

OpenStudy (misty1212):

then don't "move the 4" subtract 4

OpenStudy (misty1212):

sorry, typo there i meant \[y+4=-\frac{8}{3}x+24\] i dropped the x by mistake

OpenStudy (anonymous):

wait, where did 24 come? distributing the slope on the right side?

OpenStudy (misty1212):

yeah

OpenStudy (misty1212):

cancel the 3's

OpenStudy (misty1212):

\[-\frac{8}{3}\times (-9)=3\times 8=24\]

OpenStudy (anonymous):

Okay, I understand.

OpenStudy (misty1212):

so now we are at \[y+4=-\frac{8}{3}x+24\] subtract 4 from both sides and you are done

OpenStudy (anonymous):

y=-8/3x+20

OpenStudy (misty1212):

yup

OpenStudy (misty1212):

took a while but i hope it is ok

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (anonymous):

Thanks by the way (:

OpenStudy (misty1212):

yeah it is the "final answer" as they say on tv you are quite welcome \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

haha, of course! I'll give you a medal.

OpenStudy (misty1212):

thanks!